OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f + g)(5)? 325 9750 355 295

OpenStudy (kropot72):

To solve this you need to substitute 5 for x in both equations and add the results. Do you want me to help?

OpenStudy (anonymous):

yes, please!

OpenStudy (kropot72):

Substituting 5 for x in the first equation gives: \[f(x)=13x ^{2}\] \[f(5)=13\times 5^{2}=you\ can\ calculate\]

OpenStudy (anonymous):

i got 325

OpenStudy (kropot72):

That is correct. Now for the second equation: \[g(x)=7x-5\] \[g(5)=(7\times 5)-5=you\ can\ calculate\]

OpenStudy (anonymous):

30?

OpenStudy (kropot72):

Correct! \[(f+g)(5)=325+30=?\]

OpenStudy (anonymous):

335?

OpenStudy (kropot72):

Not really. 325 + 30 = ?

OpenStudy (anonymous):

I thought it was 335

OpenStudy (kropot72):

Have you tried adding on a calculator 325 + 30?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

What was the result on the calculator?

OpenStudy (anonymous):

355. Oh sorry! I see what i did.

OpenStudy (kropot72):

Yes, the correct choice is 355.

OpenStudy (anonymous):

can u help me with 4 more similar questions?

OpenStudy (kropot72):

Sorry, I need to log out now. Please post your questions and others can help you.

OpenStudy (anonymous):

ok